OpenStudy (anonymous):

Hello, my cat has to jump to a platform at coordinate (Xa,Ya) with a jump speed of "JumpSpeed". He is currently at coordinate (0,0). And "t" = 0, when he jump. What Angle "a" should he jump ? (note It's note necessary possible) JumpVelX = JumpSpeed * cos(a) JumpVelY = JumpSpeed * sin(a) Position = t*V + (1/2)t²*g g = gravity (0,-9.81) positionX = t * JumpSpeed * cos(a) positionY = t * JumpSpeed * sin(a) + (1/2)t² * g with positionX we get : t = positionX/(cos(a)*JumpSpeed) ....

OpenStudy (anonymous):

lets replace positionY's "t" with it : positionY = t * JumpSpeed * sin(a) + (1/2)t² * g positionY = (positionX/(cos(a)*JumpSpeed)) * JumpSpeed * sin(a) + (1/2)(positionX/(cos(a)*JumpSpeed))² * g But i can't resolver "a" (angle).

OpenStudy (anonymous):

\[PositionY = \frac{ PositionX }{ \cos(a) * JumpSpeed } * JumpSpeed * \sin(a) + \frac{ 1 }{ 2 } * \frac{ PositionX }{ \cos(a) * JumpSpeed } * g\]